# Physics-Informed Inductive Biases for Voltage Prediction in Distribution Grids

### v1

@@ 1 Introduction
Distribution networks are shifting from passive delivery systems to actively managed infrastructures shaped by high DER penetration, electrification, and frequent reconfigurations. These conditions create volatile conditions and tight real-time decision windows. Voltage prediction is a central task in this setting: maintaining voltages within operational limits is essential for efficiency, equipment safety, and system stability. Moreover, in distribution grids, voltage prediction is the main target as these networks typically contain only a slack bus and load buses. For an $n$-bus distribution system, the AC power-flow problem determines bus voltage magnitudes and phase angles from specified injections based on a coupled system of 2(n1)2(n-1) nonlinear equations. Classical solvers can be accurate but scale poorly and may fail to converge under realistic operating conditions. Simplified linear methods are computationally efficient but unreliable for distribution grids, where higher resistance-to-reactance (R/X) ratios cause more voltage drops, which can lead to greater active power losses and impact voltage stability. These limitations have motivated increasing attention to machine learning approaches, which approximate the mapping from loads and network parameters to voltages without repeated numerical iterations. Recent studies, particularly those using Graph Neural Networks (GNNs), demonstrate that ML-based solvers can achieve orders-of-magnitude speedups while retaining strong accuracy and flexibility (Donon et al., 2020; Jeddi & Shafieezadeh, 2021). Unlike tabular models that must be retrained for each topology, GNNs naturally adapt to varying grid sizes and configurations with little structural modification. This adaptability makes them well-suited for distribution systems, which are characterized by heterogeneity and frequent reconfigurations. However, purely data-driven models suffer from poor generalization, especially under unseen configurations or operating conditions, a key barrier for deployment in real grids. This motivates a systematic investigation into how inductive biases can guide ML models toward more accurate, generalizable, and reliable predictions.
@@
Inductive biases are assumptions embedded in a model’s architecture, training process, or data representation that guide learning toward specific solutions. For voltage prediction, relevant inductive biases include the integration of graph-based models, physical laws (e.g., power flow equations), or operational constraints. By incorporating these biases, models can better capture the underlying structure of the problem, leading to improved performance and robustness. Among inductive biases, physics-informed approaches have received particular attention. These models embed prior knowledge of the system’s governing equations to enforce physical consistency in a supervised or unsupervised training regime. Nevertheless, most prior research has focused narrowly on equation-constrained losses, without investigating alternative ways of embedding physical knowledge. Moreover, little is known about the impact of these inductive biases on generalization performance, which is crucial for deployment in real-world grids where unseen configurations are the norm.
@@
For each inductive bias, controlled experiments are conducted across a dataset of typical distribution grid topologies, evaluating standard predictive accuracy as well as generalization performance under unseen network configurations  [^1] All experiment models and code available at: [Under review]. Due to their inherent generalization advantage, we focus on Graph Neural Network architectures in this study. By isolating and comparing the presented inductive biases rather than model complexity alone, we provide a systematic assessment of their effectiveness in guiding learning toward physically meaningful and generalizable solutions, highlighting new pathways for learning voltage prediction. Ultimately, our findings provide concrete guidance for the design of ML models that are not only fast and accurate but also reliable for safety-critical, real-world deployment in modern distribution grids. In this light, we make the following contributions:
  • We provide a systematic assessment of physics-informed inductive biases for voltage prediction in distribution grids.
  • We perform a comprehensive benchmarking of predictive accuracy and out-of-distribution generalization across heterogeneous low- and medium-voltage networks.
  • We derive actionable design guidelines identifying which inductive biases most effectively enhance reliability and scalability, learning-based voltage prediction models.
@@ 2.2 Physics-Informed Machine Learning in Power Systems
Physics-informed machine learning integrates physical laws and domain knowledge into machine learning models to enhance their performance, interpretability, and reliability (Karniadakis et al., 2021). This approach is particularly relevant in engineering domains like power systems, where physical principles govern system behavior. By embedding these principles into the learning process, this technique aims to provide more trustworthy models that can generalize better to unseen scenarios. In power systems, physics-informed neural networks (PINN) have been applied to various tasks, including state estimation, dynamics analysis, and optimal power flow (Huang & Wang, 2023). Here we review three strategies of physics-informed machine learning that are particularly relevant to our study of voltage prediction in distribution grids.
@@ 2.2.1 Power-flow-constrained Loss Function
This is one of the most common approaches to physics-informed machine learning in power systems and was highlighted as a core PINN paradigm by a recent review paper (Huang & Wang, 2023). This is typically achieved by assigning physical meanings to the variables in the output neural networks (NNs) and embedding the governing equations of these physical variables into the loss function. The loss function can thus be formally written as:
@@
Since this method applies to the training process itself, it is model-agnostic and can be adapted for arbitrary neural networks architectures. Due to the ubiquity of this approach, we only highlight the most relevant studies here. GraphNeuralSolver (Donon et al., 2020) was one of the first works to apply this technique for power flow estimation, using their own custom GNN architecture and serving as a benchmark for many future works in this direction. (Jeddi & Shafieezadeh, 2021) integrates this method into a Graph Attention Network (GAT) model in order to learn the importance of neighbouring nodes, while (Böttcher et al., 2023) adopts a randomized, recurrent GNN architecture, so that its predictor learns how to solve the power flow problem from many different starting points. All three works solve the problem in a strictly unsupervised manner, completely skipping the need for labeled training data and the $L_{\text{pred}}$ term. Other notable works include (Zamzam & Sidiropoulos, 2020) and (Habib et al., 2024) who successfully apply this technique to distribution grid state estimation using non-GNN architectures. However, many of these works focus their analysis on the transmission system context, and none of them evaluate the generalization performance across a structured suite of distribution grid configurations.
@@ 2.2.2 Complex-Valued Neural Networks
To date, there are relatively few works that explore complex-valued neural networks for power system applications, despite the fact that electrical quantities such as voltages and currents are inherently complex-valued. For distribution grid voltage prediction, this allows voltage magnitudes and angles to be modeled jointly, rather than separately as in real-valued models. An early work developed a CVNN for load flow prediction and noted improved reliability compared to a real valued neural network baseline (Ceylan et al., 2005). However, both models were very small MLPs, consisting of one hidden layer of maximum 8 neurons, and evaluation was performed on a toy 6-bus network. Thus, the results seem hardly tractable for a real-world setting. A more comprehensive study is presented in (Wu et al., 2023), which introduces a complex-valued spatio-temporal graph convolutional network (Cplx-STGCN) for the tasks of power system state forecasting and false data injection detection, and highlight the advantages of this model against several baselines. However, none of the analyzed baselines include a real-valued counterpart of Cplx-STGCN. Moreover, the second complex-valued model presented, CplxFNN, performs poorly. Additionally, the evaluation examines a single IEEE 118-bus transmission system and the topology change experiment is limited to the removal of a single line. Therefore, the results remain largely inconclusive of the direct effects of CVNNs for distribution grid voltage prediction. Though not directly a machine learning model, the recently-proposed Physics-Informed Symbolic Regression (PISR) method leverages complex-valued representations to learn analytical approximations of the power flow equations (Eichhorn et al., 2025). The model embeds this complex information directly into the symbolic regression process, drastically elevating the performance of the non-physics-constrained variant, and allowing the model to learn more reliably with less data samples compared to the MLP baseline. Similarly, this study is limited to two small-scale systems, but the results are indicative of the promise of complex-valued representations for power system tasks.
@@ 2.2.3 Residual Prediction
Due to the complex physics of power systems, machine learning training can be unstable and heavily reliant on initializations and hyperparameters. However, voltage magnitudes and angles are typically close to a constant baseline set by the slack bus, thus motivating the use of residuals. By focusing on capturing deviations, the redundant learning of the trivial baseline is eliminated thereby reducing complexity. Currently, state of the art power-flow methods typically predict the absolute voltage targets directly. Residual learning was employed for probabilistic power flow (Chen & Zhang, 2023) and simulated on standard transmission system test cases. Particularly, both model-based and purely data-driven versions of weights initialization were utilized for the residual layer.
@@ 3.1 Experiment Setup
For controlled comparison of the three inductive biases, we conduct a series of experiments using a consistent dataset, model architecture, and evaluation metrics. The ENGAGE dataset (Okoyomon, 2025) is a collection of low and medium voltage grids based on the SimBench networks (Meinecke et al., 2020). This test suite was introduced by Okoyomon & Goebel (2025) to evaluate the generalization capabilities of power flow models across several distribution grid configurations and promote robust grid planning and operation. With 3000 sample networks, derived based on structural and data variations of 10 base distribution grids, the dataset contains test cases with varying sizes, topologies, and electrical characteristics, making it ideal for evaluating generalization performance. We first establish a baseline to identify how well each model performs under known network configurations and learned data distributions. We randomly shuffle the data with all grids and adopt a 75/15/10 train/validation/test split for model training. To evaluate generalization performance, we conduct an Out-of-Distribution (OOD) experiment. We use a leave-one-out approach where we train on all but one of the base grid configurations and test on the held-out configuration. We thus apply the same train/validation/test split, with 300 samples (10%) in the held-out configuration. This approach allows us to assess how well each model can generalize to unseen grid topologies and operating conditions, which is critical for real-world deployment.
To evaluate model performance, we employ the standard metrics of Root Mean Squared Error (RMSE) and Mean Absolute Percentage Error (MAPE) between the predicted and true voltage magnitudes and angles across all buses in the network. Additionally, we report the model training time to assess computational efficiency. To evaluate generalization performance, the RMSE of each held-out configuration is compared against the RMSE on the baseline test set. A smaller increase in RMSE/MAPE on the held-out configuration indicates better generalization.
@@ 3.2 Model Design
The model utilizes two-stage preprocessing and postprocessing layers, implemented using fully connected networks with ReLU activations and batch normalization. The core processing consists of 7 message-passing layers using the GraphConv update scheme:
$x_{i}^{{}^{\prime}}=W_{1}x_{i}+W_{2}\sum_{j\in N(i)}e_{ij}x_{j}$
For all other models, we keep the architecture as similar as possible to the baseline, only changing the components necessary to implement each inductive bias. This approach allows us to isolate the effects of each inductive bias while minimizing confounding factors. We also apply uniform training procedures to ensure a fair comparison, using a mean squared error as a supervised loss function, with 3000 epochs, a batch size of 128, and an Adam optimizer with an adaptive learning rate starting from $10^{-3}$ until $10^{-5}$. Using a hyperparameter tuning, we determine the optimal model configuration and ensure strong performance on standard accuracy metrics before proceeding with the experiment variants.
@@ 3.3 Power-flow-constrained Loss Model
We implement a loss function that combines the standard supervised loss with a physics-based regularization that penalizes violations of the AC power flow equations. More specifically, we train the baseline GNN as usual, and then fine-tune the model using the physics informed loss for 20 epochs. This scheme proved to be much more reliable than training completely unsupervised or with a weighted physics loss, since these methods are highly reliant on hyperparameters, with initial loss values beginning as high as $10^{8}$. Furthermore, since the data is derived using pandapower (Thurner et al., 2018), we replicate their AC power balance equations and model assumptions to implement our loss function. We ensure that the physics informed loss function provides near-zero values for the ground truth voltage targets before using it for model training.
@@ 3.5 Residuals Model
We implement a GNN architecture that predicts the residuals as deviations from the slack bus voltage, rather than predicting absolute voltage values directly. This approach simplifies the learning task, as the model only needs to learn the differences from a known baseline (the slack bus voltage), which are often smaller in magnitude and may exhibit simpler patterns than the original target variable. We implement the residuals model by enforcing a single skip connection in the model, that connects the neural network’s output to the slack bus reference value (Donti et al., 2017). This way, the model learns to output residuals, which are added to the slack bus reference values to produce the final voltage predictions. This allows us to keep the input data format consistent across all models, simplifying data handling and preprocessing.
@@ 4 Results
We evaluate three inductive-bias variants for voltage prediction in distribution grids compared to a purely data-driven baseline model. We report performance for predicting grids with familiar structures, and use out-of-distribution experiments to examine model generalization. The evaluation networks are SimBench low-voltage (LV) and medium-voltage (MV) systems, consisting of a mixture of rural, semiurban, urban, and commercial networks. In this section, we present general model performance results across all experiments using RMSE and time metrics. For a complete view of individual benchmarking results, we refer the reader to Appendix A.
Table 1 provides a statistical summary of the generalization experiments, while Figure 1 displays RMSE and training time results for every test case in the study. Voltage magnitude RMSE is presented using p.u. values, meaning that results are normalized with respect to the reference voltage of the network, allowing for interpretation as relative errors (e.g., 0.01 p.u. is equivalent to 1% of the nominal voltage). In terms of voltage magnitude, the supervised physics-informed (PI) variant achieves the highest reliability across the simulated topologies, indicating greater robustness as a magnitude predictor. The complex-valued model attains the best angle accuracy, with an improvement of two orders of magnitude across all experiments, at the cost of increased computation time (x4). Although the residual learning model performs the best in certain cases, we observe that it does not provide any significant advantage, and exhibits low performance consistency with higher standard deviation. Generally, both the PI and complex-valued variants are the top performers, with relative advantage depending on the specific grid under consideration.
@@
Complex 0.0102 0.0584 0.0189 0.0142 0.0093 0.0651 0.0212 0.0179
Phys-Loss 0.0070 0.0266 0.0166 0.0060 0.6013 2.3928 1.2628 0.5595
Residuals 0.0094 0.0389 0.0183 0.0083 0.6101 4.2366 1.3419 1.0598
(a) Voltage Magnitude
(b) Voltage Angle
(c) Training time
Figure 1: Summary of selected voltage prediction performance across all experiments.

### v2

@@ 1 Introduction
Distribution networks are shifting from passive delivery systems to actively managed infrastructures shaped by high DER penetration, electrification, and frequent reconfigurations. These conditions create volatile conditions and tight real-time decision windows. Voltage prediction is a central task in this setting because maintaining voltages within operational limits is essential for efficiency, equipment safety, and system stability. For an $n$-bus distribution system, typically containing only a slack bus and load buses, the AC power-flow problem determines bus voltage magnitudes and phase angles from specified injections based on a coupled system of 2(n1)2(n-1) nonlinear equations. Classical solvers can be accurate but scale poorly and may fail to converge under realistic operating conditions. Simplified linear methods are computationally efficient but unreliable for distribution grids, where higher resistance-to-reactance (R/X) ratios cause more voltage drops, which can lead to greater active power losses and impact voltage stability. These limitations have motivated increasing attention to machine learning approaches, which approximate the mapping from loads and network parameters to voltages without repeated numerical iterations. Recent studies, particularly those using Graph Neural Networks (GNNs), demonstrate that ML-based solvers can achieve orders-of-magnitude speedups while retaining strong accuracy and flexibility (Donon et al., 2020; Jeddi & Shafieezadeh, 2021). Unlike tabular models that must be retrained for each topology, GNNs naturally adapt to varying grid sizes and configurations with little structural modification. This adaptability makes them well-suited for distribution systems, which are characterized by heterogeneity and frequent reconfigurations. However, purely data-driven models suffer from poor generalization, especially under unseen configurations or operating conditions, a key barrier for deployment in real grids. This motivates a systematic investigation into how inductive biases can guide ML models toward more accurate, generalizable, and reliable predictions.
@@
Inductive biases are assumptions embedded in a model’s architecture, training process, or data representation that guide learning toward specific solutions. For voltage prediction, relevant inductive biases include the integration of graph-based models, physical laws, or operational constraints. These biases help models better capture the underlying structure of the problem, leading to improved performance and robustness. Among inductive biases, physics-informed approaches have received particular attention as they embed prior knowledge of the system’s governing equations to enforce physical consistency in a supervised or unsupervised training regime. Nevertheless, most prior research has focused narrowly on equation-constrained losses, without investigating alternative ways of embedding physical knowledge. Moreover, little is known about the empirical impact of these inductive biases on generalization performance, which is crucial for deployment in real-world grids where unseen configurations are the norm.
@@
For each inductive bias, controlled experiments are conducted across a dataset of common distribution grid topologies, evaluating standard predictive accuracy as well as generalization performance under unseen network configurations  [^1] All experiment models and code available at: [Under review]. Due to their inherent generalization advantage, we focus on Graph Neural Network architectures in this study. By isolating and comparing the presented inductive biases, we provide a systematic assessment of their effectiveness in guiding learning toward physically meaningful and generalizable solutions, highlighting new pathways for learning voltage prediction. Ultimately, our findings provide concrete guidance for the design of ML models that are not only fast and accurate but also reliable for safety-critical, real-world deployment in modern distribution grids. In this light, we make the following contributions:
  • We provide a systematic assessment of standard accuracy and out-of-distribution generalization of physics-informed inductive biases for voltage prediction across heterogeneous distribution networks.
  • We contribute the first complex-valued neural network architecture for distribution grid voltage prediction and demonstrate its generalization advantage.
  • We derive actionable design guidelines, identifying which methods most effectively enhance real-world reliability and where more research is required.
@@ 2.2 Physics-Informed Machine Learning in Power Systems
Physics-informed (PI) machine learning integrates physical laws and domain knowledge into machine learning models to enhance their performance, interpretability, and reliability (Karniadakis et al., 2021). This approach is particularly relevant in engineering domains like power systems, where physical principles govern system behavior. By embedding these principles into the learning process, this technique aims to provide more trustworthy models that can generalize better to unseen scenarios. In power systems, physics-informed neural networks (PINN) have been applied to various tasks, including state estimation, dynamics analysis, and optimal power flow (Huang & Wang, 2023). Although these approaches claim to improve model generalization, there have been no dedicated generalization studies of these methods for power systems, thus motivating our systematic assessment. In the following, we review three strategies of physics-informed machine learning that are particularly relevant to our study of voltage prediction in distribution grids.
@@ 2.2.1 Power-flow-constrained Loss Function
Constraining loss functions is one of the most common approaches to physics-informed machine learning in power systems and was highlighted as a core PINN paradigm by a recent review paper (Huang & Wang, 2023). This is typically achieved by assigning physical meanings to the variables in the output neural networks (NNs) and embedding the governing equations of these physical variables as an additional loss term. The loss function can thus be formally written as:
@@
Since this method applies to the training process itself, it is model-agnostic and can be adapted for arbitrary neural network architectures. Due to the ubiquity of this approach, we only highlight the most relevant studies here. GraphNeuralSolver (Donon et al., 2020) was one of the first works to apply this technique for power flow estimation, using their own custom GNN architecture and serving as a benchmark for many future works in this direction. (Jeddi & Shafieezadeh, 2021) integrates this method into a Graph Attention Network (GAT) model in order to learn the importance of neighbouring nodes, while (Böttcher et al., 2023) adopts a randomized, recurrent GNN architecture, so that its predictor learns how to solve the power flow problem from many different starting points. All three works solve the problem in a strictly unsupervised manner, completely skipping the need for labeled training data and the $L_{\text{pred}}$ term. Other notable works include (Zamzam & Sidiropoulos, 2020) and (Habib et al., 2024) who successfully apply this technique to distribution grid state estimation using non-GNN architectures. However, many of these works focus their analysis on the transmission system context, and none of them evaluate the generalization performance across a structured suite of distribution grid configurations.
@@ 2.2.2 Complex-Valued Neural Networks
To date, there are relatively few works that explore complex-valued neural networks for power system applications, despite the fact that electrical quantities such as voltages and currents are inherently complex-valued. For distribution grid voltage prediction, this allows voltage magnitudes and angles to be modeled jointly, rather than separately as in real-valued models. Though not directly a neural network model, the recently-proposed Physics-Informed Symbolic Regression (PISR) method leverages complex-valued representations to learn analytical approximations of the power flow equations (Eichhorn et al., 2025). The model embeds complex information directly into the symbolic regression process, drastically elevating the performance and allowing the model to learn more reliably with less data samples compared to the MLP baseline. This study is limited to two small-scale systems, but the results are indicative of the promise of complex-valued representations for power system tasks. A much earlier work made a first attempt at a CVNN for load flow prediction and noted improved reliability compared to a real valued neural network baseline (Ceylan et al., 2005). However, both models were very small MLPs, consisting of one hidden layer of maximum 8 neurons, and evaluation was performed on a toy 6-bus network. Thus, the results seem hardly tractable for a real-world setting. Lastly, a more comprehensive study is presented in (Wu et al., 2023), which introduces a complex-valued spatio-temporal graph convolutional network (Cplx-STGCN) for the task of power system state forecasting (PSSE) and highlights the advantages of this model against several baselines. However, their performance increase cannot be attributed to the complex valued network, as their CplxFNN model performs poorly and they do not evaluate real-valued counterpart of Cplx-STGCN. Similarly, their evaluation is limited to a single IEEE Additionally, the evaluation examines a single IEEE 118-bus transmission system. Furthermore, the PSSE task is fundamentally different from our setting as it aims to predict unknown voltages in a network using historical voltages and a set of available measurements, while our voltage prediction task aims to estimate complex nodal voltages given only complex power measurements, effectively learning the power balance equations. To our knowledge, this study contributes the first evaluation of complex valued neural networks for voltage prediction in a realistic distribution grid setting and illustrates their efficacy in robust voltage estimation for out-of-distribution use cases.
@@ 2.2.3 Residual Prediction
Due to the complex physics of power systems, machine learning training can be unstable and heavily reliant on initializations and hyperparameters. Currently, state of the art power-flow methods typically predict the absolute voltage targets directly. However, voltage magnitudes and angles are generally close to a constant baseline set by the slack bus, thus motivating the use of residuals. By focusing on capturing deviations, the redundant learning of the trivial baseline is eliminated, thereby reducing complexity. Residual learning was employed in (Chen & Zhang, 2023) for probabilistic power flow and simulated on standard transmission system test cases. Beyond voltage prediction, (Eddin Za’ter et al., 2025) use residual learning to learn the nonlinear residual between a given DC-OPF solution and the corresponding AC-OPF solution. When compared against models that directly learn the AC-OPF solution, the residual-learning framework achieved a 35–45% reduction in voltage, power, and feasibility errors, demonstrating its effectiveness for OPF prediction. The success of this approach suggests that residual learning may also be advantageous for voltage prediction.
@@ 3.1 Experiment Setup
The experimental overview is presented in Figure 1. We evaluate a baseline model and physics-informed variants in the tasks of in-distribution and out-of-distribution voltage prediction. For controlled comparison of the three inductive biases, we conduct a series of experiments using a consistent dataset, model architecture, and evaluation metrics. The ENGAGE dataset is a collection of low and medium voltage grids based on the SimBench networks. This test suite was introduced by Okoyomon & Goebel (2025) to evaluate generalization capabilities of power flow models across several distribution grid configurations and promote robust grid planning and operation. With 3000 sample networks derived from 10 base distribution grids, the dataset contains test cases with varying sizes, topologies, and electrical characteristics, making it ideal for evaluating generalization performance. A summary of the experiment data and task formulation can be found in Appendix C.
We first establish a baseline to identify how well each model performs under known network configurations and learned data distributions. We randomly shuffle the data with all grids and adopt a 75/15/10 train/validation/test split for model training. To evaluate generalization performance, we conduct an out-of-distribution (OOD) experiment. We use a leave-one-out approach where we train on all but one of the base grid configurations and test on the held-out configuration. We thus apply the same train/validation/test split, with 300 samples (10%) in the held-out configuration. This approach allows us to assess how well each model can generalize to unseen grid topologies and operating conditions, which is critical for real-world deployment. Though a one-to-one grid evaluation is possible, previous studies indicate that current ML solvers are not mature enough for this use case (Böttcher et al., 2023; Hansen et al., 2022; Jeddi & Shafieezadeh, 2021; Okoyomon & Goebel, 2025). As such, we focus on the OOD setting for our generalization study.
To evaluate model performance, we employ the standard metric of Root Mean Squared Error (RMSE) between the predicted and true voltage magnitudes and angles across all buses in the network. Additionally, we report the training time, inference time, and model capacity to assess computational efficiency. To evaluate generalization performance, the RMSE of each held-out configuration is compared against the RMSE on the baseline test set. A tighter and lower RMSE distribution on the held-out configuration indicates better generalization.
Figure 1: Experiment methodology for in-distribution and out-of-distribution model evaluation
@@ 3.2 Model Design
The model utilizes two-stage preprocessing and postprocessing layers, implemented using fully connected networks with ReLU activations and batch normalization. The core processing consists of 7 128-dimensional message-passing layers using the GraphConv update scheme. To attribute generalization performance to inductive bias rather than the specific GNN backbone, we validate results through additional experimentation on a GAT backbone (Brody et al., 2021) in Appendix B.2.
For all models, we keep the architecture as similar as possible to the baseline, only changing the components necessary to implement each inductive bias. This allows us to isolate the effects of each inductive bias while minimizing confounding factors. As an additional analysis, we present results of combined inductive bias models in Appendix B.1. For all models, we apply uniform training procedures to ensure a fair comparison, using MSE as the supervised loss function, 3000 epochs, batch sizes of 128, and an Adam optimizer with an adaptive learning rate starting from $10^{-3}$ until $10^{-5}$. Using a hyperparameter tuning, we determine the optimal model configuration and ensure strong performance on standard accuracy metrics before proceeding with the experiment variants.
@@ 3.3 Power-flow-constrained Loss Model
We implement a loss function that combines the standard supervised loss with a physics-based regularization that penalizes violations of the AC power flow equations. More specifically, we train the baseline GNN as usual, and then fine-tune the model using the physics informed loss for 20 epochs. This limit was deliberately chosen to ensure this phase acts solely as a structural refinement, preventing it from becoming a deep re-optimization of the landscape. We observed no significant gain in OOD performance when extending this phase up to 30 epochs. This scheme proved to be much more reliable than training completely unsupervised or with a weighted physics loss, since these methods are highly reliant on hyperparameters. Furthermore, since the data is derived using pandapower (Thurner et al., 2018), we replicate their AC power balance equations and model assumptions to implement our loss function. We ensure that the physics informed loss function provides near-zero values for the ground truth voltage targets before using it for model training.
@@ 3.5 Residuals Model
We implement a GNN architecture that predicts the residuals as deviations from the slack bus voltage. This simplifies the learning task, as the model only needs to learn the differences from a known baseline (the slack bus voltage), which are often smaller in magnitude and may exhibit simpler patterns than the original target variable. We implement the residuals model by enforcing a single skip connection that connects the neural network’s output to the slack bus reference value (Donti et al., 2017). This way, the model learns to output residuals, which are added to the slack bus reference values to produce the final voltage predictions. This allows us to keep the input data format consistent across all models, simplifying data handling and preprocessing.
@@ 4 Results
We evaluate three inductive-biases for voltage prediction in distribution grids compared to a purely data-driven baseline model. We report performance for prediction on grids with familiar structures and on unseen (i.e. OOD) grids. In this section, we present general model performance results using predictive performance and model efficiency metrics. For a complete view of individual benchmarking results and additional experimentation, we refer the reader to Appendices A and B.
Table 1 provides a statistical summary of the generalization experiments, while Figure 2 displays overall performance metrics. Voltage magnitude RMSE is presented using p.u. values, meaning that results are normalized with respect to the reference voltage of the network, allowing for interpretation as relative errors (e.g., 0.01 p.u. is equivalent to 1% of the nominal voltage).
In terms of voltage magnitude, the supervised physics-informed variant achieves the highest reliability across the simulated topologies, indicating greater robustness as a magnitude predictor. The complex-valued model attains the best angle accuracy, with an improvement of two orders of magnitude across all experiments, at the cost of extra computational burden. Although the residual learning model performs the best in certain cases, it does not provide any significant advantages.
@@
Res 0.0094 0.0389 0.0183 0.0083 0.6101 4.2366 1.3419 1.0598
Cplx 0.0102 0.5842 0.0715 0.1802 0.0093 0.0651 0.0212 0.0179
PFLoss 0.0070 0.0266 0.0166 0.0060 0.6013 2.3928 1.2628 0.5595